# Beyond Gaze Points: Augmenting Eye Movement with Brainwave Data for Multimodal User Authentication in Extended Reality

### v1

@@ Abstract.
The increasing adoption of Extended Reality (XR) in various applications underscores the need for secure and user-friendly authentication methods. However, existing methods can disrupt the immersive experience in XR settings, or suffer from higher false acceptance rates. In this paper, we introduce a multimodal biometric authentication system that combines eye movement and brainwave patterns, as captured by consumer-grade low-fidelity sensors. Our multimodal authentication exploits the non-invasive and hands-free properties of eye movement and brainwaves to provide a seamless XR user experience and enhanced security as well. Using synchronized eye and brainwave data collected from 30 participants through consumer-grade devices, we investigated whether twin neural networks can utilize these biometrics for identity verification. Our multimodal authentication system yields an excellent Equal Error Rate (EER) of 0.298%, which means an 83.6% reduction in EER compared to the single eye movement modality or a 93.9% reduction in EER compared to the single brainwave modality.
@@ 1. Introduction
To provide empirical validation for our multimodal authentication approach in XR, we conducted a lab study with 30 participants, using consumer-grade equipment to record both brainwaves and eye movements (Sec 3). Leveraging this synchronized dataset, we develop an authentication system employing twin neural networks based (Sec 4) on an interactive dot stimulus (Sluganovic et al., 2018). We implemented various feature fusion and score fusion strategies to investigate the most promising configurations. Additionally, since the pupil diameter is used as a feature by some papers (Eberz et al., 2019; Zhang et al., 2018), but others omit it (Sluganovic et al., 2018; Lohr and Komogortsev, 2022); We investigate the effect of pupil diameter on authentication performance to illuminate its exact influence on the system. The decision to include pupil diameter in feature engineering may be influenced by the designer’s choice or the unavailability of pupil diameter. Our results demonstrate a notably low Equal Error Rate (EER): 0.686% without pupil diameter and 0.298% with pupil diameter using score fusion (Sec 5). Before concluding (Sec 9), we discuss the technical feasibility of XR integration (Sec 6), compare our results with related work (Sec 7), and also discuss the limitations of our work (Sec LABEL:Section:limit). Accordingly, we offer the following contributions:
@@
2. Background
We begin by defining authentication. Next, we introduce our adversary model, articulating the specific problem we aim to solve in a ‘zero-effort attacker’ scenario. Following that, we discuss state-of-the-art authentication, focusing on XR authentication. The discussion then advances to explain the utilization of neural network architectures in biometric authentication, detailing how they contribute to enhancing the system’s accuracy.
2.1. Authentication
Biometric recognition systems commonly perform either of two fundamental tasks: verification or identification, as standardized by the ISO international standard (ISO/IEC 2382-37:2022, en). Identity verification, also known as authentication, implements a 1-1 comparison to confirm whether an individual presenting a previously registered biometric is indeed the same person. Essentially, it answers the question, ”Are you who you claim to be?” This is primarily utilized in user authentication scenarios to distinguish between legitimate users and fraudulent actors. Identification involves a one-to-many comparison with the aim of determining if the presented biometric matches any specific individual in the system’s database, essentially answering the question, ”Who are you?”.
Biometric authentication consists of two main phases: enrollment and verification. In the enrollment phase, users undergo data collection, where their unique biometric traits are captured and stored as a template $t\in T$. Each of the templates is then associated with a distinct user identifier $i\in I$, such as a username. During the verification phase, the system compares the presented biometric sample $s_{c}$ with the stored template $t_{i}$ that corresponds to the claimed identifier $i_{{c}}$. Access to the system is either granted or denied based on this comparison, formalized by the following hypotheses:
$H_{0}:M(F(s_{c}),t_{i_{{c}}})\geq\theta\quad\text{vs.}\quad H_{1}:M(F(s_{c}),t_{i_{{c}}})<\theta$
In this formalization, $F$ is the feature extraction function that takes as an input a biometric sample and outputs a vector from the template space $T$, $M$ is a matching function that outputs a distance between two inputs from the template space, and $\theta$ is a threshold for acceptance. The null hypothesis $H_{0}$ posits that the presented biometric sample does not match the stored template corresponding to the claimed identifier, in other words rejecting the request, while the alternative hypothesis $H_{1}$ posits that the sample is considered to match the template.
2.2. Adversary Model
We aim to develop a system tailored for authenticating users in either XR or desktop applications for secondary security measures. In the context of this work, we consider zero-effort attacks (Mansfield and Wayman, 2002) as the primary adversary model. The zero-effort attack in biometric authentication refers to an attack where the adversary presents their own biometric sample ($s_{c}$) to gain unauthorized access to another individual’s account ($i_{\text{c}}$). In this context, the security objective of the biometric system is to minimize the likelihood that the authentication algorithm incorrectly verifies the adversary’s identity as the legitimate user (H1). Specifically, the probability of successful verification of the wrongly claimed identity, based on the sensor readings presented by the adversary, should be negligible.
2.3. Authentication in XR
@@
Authentication methods in extended reality (XR) can be broadly classified into three primary categories: Knowledge-Based methods (Mathis et al., 2020; Düzgün et al., 2022), Possession-Factors methods (Chan et al., 2015), and Inherence-Factors methods (biometrics)(Eberz et al., 2019; Li et al., 2016; Boutros et al., 2020). Knowledge-Based methods involve what the user knows, like passwords or PINs. Possession-Factors utilize physical objects or tokens that the user owns, such as mobile phones or hardware dongles. Inherence-Factors, or biometrics, exploit unique physiological or behavioral traits for authentication.
@@
In XR settings, authentication based on knowledge and possession factors are subject to usability challenges. Knowledge-Based methods often involve credentials that are difficult to remember and also difficult to enter, while authentication based on possession factors require the availability of a physical token, the corresponding port on the device, and could be subject to theft (Stephenson et al., 2022). Biometric authentication eliminates the need to memorize passwords or to carry external tokens for access. Further, biometrics have become a viable option for authentication in XR settings because new sensors are continually being integrated into XR devices to enhance the immersive experience for users (Stephenson et al., 2022).
@@
Authentication based on inherence (biometric) factors, for instance using eye movement and brainwaves, have already gained attention beyond the realm of XR (Arias-Cabarcos et al., 2023; Sluganovic et al., 2018). For example, authentication by eye movement and brainwave is both hands-free and difficult to spoof. Consequently, Inherence-Factor (biometric) methods offer a more user-friendly alternative. Still, these methods come with their own sets of challenges. To operate effectively, a biometric system requires the trait used in authentication to be universal (i.e., present in all individuals) and permanent (i.e., stable over time). Furthermore, the trait for authentication must also satisfy a whole line of further requirements. The trait must be unique and quantifiable, and it must also be resistant to forgery. The biometric trait must be revocable by the user, who will need to consider, as well, that authentication by that trait is acceptable. The price for sensing the trait must be affordable, and the performance of the sensing must be reliable.
@@
Studies have demonstrated the temporal stability of brainwave and eye movement-based biometric authentication methods. Maiorana (Maiorana, 2021) conducted a year-long study, confirming the viability of brainwave authentication. Similarly, Lohr et al.(Lohr and Komogortsev, 2022) established that eye movement based authentication remains effective over a period of three years. While static biometrics, such as fingerprints or iris scans, cannot be altered if compromised. Lin et al. (Lin et al., 2018) provided empirical data suggesting that brainwave passwords can be modified by assigning a different task. In terms of usability, several studies have indicated a general willingness among people to adopt eye movement and brainwave authentication methods, although privacy concerns still remain to be addressed (Fallahi et al., 2024; Röse et al., 2023; Brooks et al., 2013). Overall, we acknowledge the need for further research to examine other aspects of biometric authentication, but the specific aim of our work is to achieve reliable performance using sensors compatible with XR devices.
@@
2.4. Eye movement and Brainwaves
In recent years, eye movements and brainwaves have gained considerable interest as biometrics, primarily due to rapid advancements in sensor technologies that capture these physiological features. The literature demonstrates a correlation between brainwaves (EEG) and eye movements (Keren et al., 2010; Papadopoulou et al., 2010), establishing a basis for our hypothesis that combining these two modalities could significantly enhance biometric authentication systems. The following provides a brief background on each of them:
Eye movement: The human eye exhibits complex movements governed by six muscles, with primary types categorized as saccades and fixations (Eberz et al., 2019). Fixations involve holding the eye steady, focusing on an object to gain detailed visual information, while saccades are quick eye movements that reorient the gaze from one fixation point to another (Holland and Komogortsev, 2013). Furthermore, eye saccades are of two types: reflexive and voluntary. Reflexive saccades are automatic movements in response to visual changes, like a moving object. Voluntary saccades are deliberate, allowing you to shift focus intentionally (Walker et al., 2000). Both are crucial for interacting with your environment. Additionally, the pupil diameter is a unique feature influenced by various factors, including physiology, gender, and age (MacLachlan and Howland, 2002). Given the unique patterns in eye movements and pupil diameter, these features have been increasingly used for biometric authentication in recent years (Eberz et al., 2019; Sluganovic et al., 2018).
Brainwaves: Brainwaves or electroencephalogram (EEG) are the electrical activity originating from the firing of neurons in the human brain, and is typically measured at the scalp using specialized sensors (Niedermeyer and da Silva, 2005; Hu and Zhang, 2019). Initially developed for medical applications by (Haas, 2003) like the diagnosis of neurological disorders (Adeli and Ghosh-Dastidar, 2010), EEG technology has broadened its scope over the years. Beyond its medical applications, EEG technology has evolved to be in Brain-Computer Interfaces (BCI), enabling direct communication between the human brain and external devices (Kawala-Sterniuk et al., 2021). More recently, EEG has garnered attention in the field of biometric user recognition due to the uniqueness and complexity of brain wave patterns (Gui et al., 2019).
2.5. Twin Neural Network
A Twin Neural Network (TNN) is a specialized neural network architecture that consists of two or more identical sub-networks connected in parallel. These parallel branches are designed to produce feature embeddings, vectors containing a reduced-dimensional input representation. The architecture leverages these embeddings to assess the similarity of the inputs by comparing their feature vectors. TNN have gained attention in recent years, particularly for their efficacy in learning from limited data (Li et al., 2022).
Employing triplet loss for TNN has become popular with FaceNet, a face recognition pipeline introduced by Google researchers in 2015 (Schroff et al., 2015). The training process involves three types of inputs: an anchor, a positive sample (data from a class member), and a negative sample (data from a member of a different class). The process ensures that the embeddings for samples of the same identity have small distances, while those of distinct identities have large distances. Consequently, a similarity metric can be employed for verification or identification. Building upon the principles of FaceNet in other applications, BrainNet (Fallahi et al., 2023) has also employed TNN with triplet loss to embed high-dimensional and noisy samples of the brainwaves into a latent space for user recognition.
For the purpose of our research, the TNN architecture is particularly appealing because it can be trained with limited biometric data and still generalize to other subjects not included in the training set. Moreover, its ability to reduce dimensionality and preserve similarity in latent space makes it suitable for our authentication systems(Schroff et al., 2015; Fallahi et al., 2023).
@@ 3. Experimental Design and Procedures
Figure 1. Synchronized collection of eye movment and brain activity. Lab Streaming Layer (LSL) receives data from PsychoPy, an eye-tracker, and a brainwave device. Also, PsychoPy and the eye-tracker coordinate to execute focus tasks, updating the display and sending focus timestamps to LSL
@@ 3.1. Technological and Methodological Blueprint
Authentication Task - Reflexive Saccadic Responses: Drawing inspiration from the methodology proposed by Sluganovic et al. (Sluganovic et al., 2018), we prioritized the measurement of reflexive saccades. Their inherent stability and low susceptibility to transient cognitive states made them more reliable than voluntary saccades. Our stimulus involved displaying a dot on the screen. When the participant’s gaze fixated on this dot, it would disappear, and a new dot would appear in a random position. This process was repeated 25 times for each round (Sluganovic et al., 2018), with participants undergoing 36 rounds. We determined the number of rounds according to the experiment’s maximum time limit of 25 minutes. To prevent fatigue and maintain the participants’ concentration, we implemented a 15-second rest interval between each round.
@@ 4.1. Authentication Approach Overview
    Comparison Module: The extracted features are then compared to stored biometric templates. We used Euclidean distance ($\sqrt{\sum_{i=1}^{n}(e_{i}-v_{i})^{2}}$) to compute the distance between enrollment and verification samples. Then it has been multiplied by -1, and the final similarity score is determined. Finally, a threshold should be set for deciding whether to accept or deny authentication requests..
@@ 4.2. Preprocessing
    Standardization: We needed to have fix sample size as input our recognition system. Therefor, through resampling, we achieved a consistent count of 256 data points per second across the dataset for both of eye-movement and brainwave data.
@@
  5. (5)
    Additional Considerations: Along with primary sample data, we generated User ID, Task ID, and Round ID for each sample. Since samples within a round may be too similar, the round ID was saved to prevent comparing the same data round in both enrollment and verification phases to enhance the evaluation’s accuracy and validity.
4.3. Recognition Network: Twin Neural Network
In order to authenticate subjects, we need to extract essential information unique to each individual from biometric signals, which often contain noise. We used twin neural network with triplet loss function as the core of the feature extraction module. The triplet loss function $L$ is expressed as the Euclidean distance:
@@ 4.3. Recognition Network: Twin Neural Network
The training process, guided by the triplet loss function, refines the weights and parameters of the CNN within the Siamese sub-networks. This enables the network to produce vector templates suitable for verification or identification. Inspired by BrainNet (Fallahi et al., 2023), which achieved remarkable results with similar time-series-based problems, we employed a CNN architecture. In our exploratory phase, we also tested Long Short-Term Memory (LSTM) networks, and fully connected networks, but these approaches did not yield satisfactory results. The chosen CNN design effectively addresses the challenges posed by the high dimensionality of the data, reducing it to a manageable size while preserving the unique characteristics of individual samples.
@@ 4.5. Fusion
FRR at FAR = 1% FRR at FAR = 0.1% FRR at FAR = 0.0%
Approach Biometric S 1 S 2 S 3 S 1 S 2 S 3 S 1 S 2 S 3
Single Biometric Brainwaves 59.47 28.86 21.45 83.64 61.22 49.09 99.98 91.25 67.99
Eye-tracking with pupil 68.73 8.149 68.52 88.06 23.25 16.44 96.64 45.11 25.00
Eye-tracking without pupil 52.85 17.10 17.97 74.89 44.86 38.81 99.85 72.20 52.80
Score Fusion Mean with pupil 25.29 0.346 0.356 45.15 2.096 1.611 87.94 8.947 3.789
Product with pupil 21.88 0.236 0.239 42.96 1.794 1.226 88.46 10.09 3.794
Mean without pupil 35.39 0.990 0.873 59.03 5.857 5.442 96.91 23.09 11.61
Product without pupil 35.27 1.123 0.965 59.45 6.850 5.956 97.25 23.72 12.36
Feature Fusion Architecture A with pupil 26.29 1.066 1.101 48.99 6.062 5.262 97.79 27.28 14.69
Architecture A without pupil 41.16 3.549 3.008 65.98 15.58 10.80 95.77 39.75 20.74
@@ 5.1. Testbed and Evaluation Metrics
Metrics: The Equal Error Rate (EER) served as a summary metric, indicating the point where the False Acceptance Rate (FAR) and the False Rejection Rate (FRR) are equal. Additionally, we report FRR at specific FAR thresholds of 1%, 0.1%, and 0%. The FAR represents the success rate of an attacker in a zero-effort attack, and the goal is to achieve a lower FAR while maintaining a reasonable FRR. Moreover, high FRR may lead to additional verification attempts, which can harm the device’s usability. Therefore, a balance must be maintained between these two metrics to ensure robust security without compromising the user experience.

### v2

@@ Abstract.
Extended Reality (XR) technologies are becoming integral to daily life. However, password-based authentication in XR disrupts immersion due to poor usability, as entering credentials with XR controllers is cumbersome and error-prone. This leads users to choose weaker passwords, compromising security. To improve both usability and security, we introduce a multimodal biometric authentication system that combines eye movements and brainwave patterns using consumer-grade sensors that can be integrated into XR devices. Our prototype, developed and evaluated with 30 participants, achieves an Equal Error Rate (EER) of 0.29%, outperforming eye movement (1.82%) and brainwave (4.92%) modalities alone, as well as state-of-the-art biometric alternatives (EERs between 2.5% and 7%). Furthermore, this system enables seamless authentication through visual stimuli without complex interaction.
@@ 1. Introduction
To provide empirical validation for our multimodal authentication approach in XR, we conducted a lab study with 30 participants, using consumer-grade equipment to record both brainwaves and eye movements (Sec 3). Leveraging this synchronized dataset, we develop an authentication system employing twin neural networks based (Sec 4) on an interactive dot stimulus (Sluganovic et al., 2018). We implemented various feature fusion and score fusion strategies to investigate the most promising configurations. Additionally, since the pupil diameter is used as a feature by some papers (Eberz et al., 2019; Zhang et al., 2018), but others omit it (Sluganovic et al., 2018; Lohr and Komogortsev, 2022); We investigate the effect of pupil diameter on authentication performance to illuminate its exact influence on the system. The decision to include pupil diameter in feature engineering may be influenced by the designer’s choice or the unavailability of pupil diameter. Our results demonstrate a notably low Equal Error Rate (EER): 0.686% without pupil diameter and 0.298% with pupil diameter using score fusion (Sec 5). Before concluding (Sec 9), we discuss the technical feasibility of XR integration (Sec 6), compare our results with related work (Sec 7.4), and also discuss the limitations of our work (Sec 8). Accordingly, we offer the following contributions:
@@
2. Authentication in XR
@@
Identity verification, or authentication, performs a 1:1 comparison to confirm if a presented biometric matches the registered individual, answering the question, ”Are you who you claim to be?” It is mainly used to differentiate legitimate users from attacker. Authentication methods in XR can be broadly classified into three primary categories: Knowledge-Based methods (Mathis et al., 2020; Düzgün et al., 2022), Possession-Factors methods (Chan et al., 2015), and Inherence-Factors methods (biometrics) (Eberz et al., 2019; Li et al., 2016; Boutros et al., 2020). Knowledge-Based methods involve what the user knows, like passwords or PINs. Possession-Factors utilize physical objects or tokens that the user owns, such as mobile phones or hardware dongles. Inherence-Factors, or biometrics, exploit unique physiological or behavioral traits for authentication.
@@
In XR settings, authentication based on knowledge and possession factors are subject to usability challenges. Knowledge-Based methods often involve credentials that are difficult to remember and also difficult to enter, while authentication based on possession factors requires the availability of a physical token, the corresponding port on the device, and could be subject to theft (Stephenson et al., 2022). Biometric authentication eliminates the need to memorize passwords or to carry external tokens for access. Further, biometrics have become a viable option for authentication in XR settings because new sensors are continually being integrated into XR devices to enhance the immersive experience for users (Stephenson et al., 2022).
@@
While biometrics offer significant advantages in terms of usability, there are notable performance and accuracy issues that need to be addressed. Multimodal biometric authentication presents a potential solution to these challenges. Recently, eye movement and brainwave-based authentication have gained significant attention due to their hands-free operation, the difficulty of spoofing, and application beyond authentication (Arias-Cabarcos et al., 2023; Sluganovic et al., 2018). Therefore, combining eye movement and brainwave modalities could potentially mitigate accuracy challenges while maintaining their inherent advantages. Nevertheless, a biometric system requires the trait used in authentication to be universal (i.e., present in all individuals) and permanent (i.e., stable over time). Furthermore, the trait for authentication must also satisfy a whole line of further requirements. The trait must be unique and quantifiable, and it must also be resistant to forgery. The biometric trait must be revocable by the user, who will need to consider, as well, that authentication by that trait is acceptable. The price for sensing the trait must be affordable, and the performance of the sensing must be reliable.
@@
Studies have demonstrated the temporal stability of brainwave and eye movement-based biometric authentication methods. Maiorana (Maiorana, 2021) conducted a year-long study, confirming the viability of brainwave authentication. Similarly, Lohr et al.(Lohr and Komogortsev, 2022) established that eye movement based authentication remains effective over a period of three years. While static biometrics, such as fingerprints or iris scans, cannot be altered if compromised. Lin et al. (Lin et al., 2018a) provided empirical data suggesting that brainwave passwords can be modified by assigning a different task. In terms of usability, several studies have indicated a general willingness among people to adopt eye movement and brainwave authentication methods, although privacy concerns still remain to be addressed (Fallahi et al., 2024; Röse et al., 2023; Brooks et al., 2013). Overall, we acknowledge the need for further research to examine other aspects of biometric authentication, but the specific aim of our work is to achieve reliable performance using sensors compatible with XR devices.
@@ 3. Experimental Design and Procedures
Figure 1. Synchronized collection of eye-movment and brain activity. Lab Streaming Layer (LSL) receives data from PsychoPy, an eye-tracker, and a brainwave device. Also, PsychoPy and the eye-tracker coordinate to execute focus tasks, updating the display and sending focus timestamps to LSL
@@ 3.1. Technological and Methodological Blueprint
Authentication Task - Reflexive Saccadic Responses: Building on the methodology proposed by Sluganovic et al.(Sluganovic et al., 2018), we focused on measuring reflexive saccades due to their inherent stability and low susceptibility to temporary changes in mental or emotional conditions, such as attention, mood, or stress—referred to as transient cognitive states. Unlike voluntary saccades, reflexive saccades are driven by automatic mechanisms, making them more reliable for authentication. Our stimulus involved presenting a dot on the screen. When the participant’s gaze fixated on the dot, it would disappear, and a new dot would appear in a random position. This sequence was repeated 25 times per round (Sluganovic et al., 2018), with participants completing 36 rounds. The number of rounds was determined based on the experiment’s 25-minute time limit. To prevent fatigue and maintain focus, we included a 15-second rest interval between rounds. Previous studies by Sluganovic et al.(Sluganovic et al., 2018) have demonstrated that this interactive dot task resists replay attacks effectively, as participants must respond to new random dot positions in real-time.
@@ 4.1. Authentication Approach Overview
    Comparison Module: The extracted features are then compared to stored biometric templates. We used Euclidean distance ($\sqrt{\sum_{i=1}^{n}(e_{i}-v_{i})^{2}}$) to compute the distance between enrollment and verification samples. Then it has been multiplied by -1, and the final similarity score is determined. Finally, a threshold should be set for deciding whether to accept or deny authentication requests.
@@ 4.2. Preprocessing
    Standardization: We needed to have a fixed sample size as input for our recognition system. Therefore, through resampling, we achieved a consistent count of 256 data points per second across the dataset for both of eye-movement and brainwave data.
@@
4.3. Feature extraction: Twin Neural Network
To authenticate subjects, we extract unique individual information from eye movement and brainwave signals, which often include noise. We employed a Twin Neural Network (TNN) with a triplet loss function as the core of our feature extraction module. A TNN is a specialized neural network architecture consisting of two or more identical sub-networks connected in parallel. The triplet loss function ensures that embeddings from the same identity are close, while embeddings from different identities are far apart  (Schroff et al., 2015). The triplet loss function $L$ is expressed as the Euclidean distance:
@@ 4.3. Feature extraction: Twin Neural Network
We employed a CNN for our sub-network architecture due to its demonstrated effectiveness in various brainwave and eye movement authentication studies (Fallahi et al., 2023; Lohr and Komogortsev, 2022; Maiorana, 2021; Bidgoly et al., 2022), specifically using the CNN architecture proposed in the BrainNet paper(Fallahi et al., 2023).
@@ 5.1. Testbed and Evaluation Metrics
Metrics: The Equal Error Rate (EER) served as a summary metric, indicating the point where the False Acceptance Rate (FAR) and the False Rejection Rate (FRR) are equal. Additionally, we report FRR at specific FAR thresholds of 1%, 0.1%, and 0.01%. The FAR represents the success rate of an attacker in a zero-effort attack, and the goal is to achieve a lower FAR while maintaining a reasonable FRR. Moreover, high FRR may lead to additional verification attempts, which can harm the device’s usability. Therefore, a balance must be maintained between these two metrics to ensure robust security without compromising the user experience.
The EER serves as a useful comparison metric across studies but is not directly practical for real-world applications. Notably, NIST (2023)/ISO  [^6] https://pages.nist.gov/800-63-3/sp800-63b.html and the European Border Guard Agency Frontex  [^7] https://www.frontex.europa.eu/assets/Publications/Research/Best_Practice_Technical_Guidelines_ABC.pdf specify that biometric systems must operate at FAR $\leq 0.1\%$. Meanwhile, FIDO  [^8] https://fidoalliance.org/specs/biometric/requirements/Biometrics-Requirements-v4.0.1-fd-20240522.pdf and the upcoming NIST (August 2024)/ISO standards  [^9] https://pages.nist.gov/800-63-4/sp800-63b.html recommend an even stricter FAR $\leq 0.01\%$. Additionally, FIDO and late NIST/ISO standards propose an FRR $\leq 5\%$, ensuring 19 successful logins out of 20 attempts for legitimate users.
FRR at FAR = 1% FRR at FAR = 0.1% FRR at FAR = 0.01%
Approach Biometric S 1 S 2 S 3 S 1 S 2 S 3 S 1 S 2 S 3
Single Biometric Brainwaves 59.47 28.86 21.45 83.64 61.22 49.09 94.43 86.92 67.99
Eye-tracking with pupil 52.85 8.149 6.852 74.89 23.25 16.44 86.17 42.08 25.00
Eye-tracking without pupil 68.73 17.10 17.97 88.06 44.86 38.81 96.08 66.33 52.80
Score Fusion Mean with pupil 25.29 0.346 0.356 45.15 2.096 1.611 62.71 6.214 3.789
Product with pupil 21.88 0.236 0.239 42.96 1.794 1.226 60.99 6.995 3.794
Mean without pupil 35.39 0.990 0.873 59.03 5.857 5.442 76.60 19.04 11.61
Product without pupil 35.27 1.123 0.965 59.45 6.850 5.956 77.20 18.72 12.36
Feature Fusion Architecture A with pupil 26.29 1.066 1.101 48.99 6.062 5.262 68.25 20.69 14.69
Architecture A without pupil 41.16 3.549 3.008 65.98 15.58 10.80 81.94 27.10 20.74
5.2. Threat Model
In alignment with the methodology of Zhang et al. (Zhang et al., 2024), we consider an adversary whose goal is to access sensitive personal information—such as user accounts, photos, or financial data—or to perform unauthorized actions like initiating payments or installing malware on a user’s XR device. We assume that the adversary is knowledgeable about the authentication dot task and has physical access to the user’s XR headset. Given these assumptions and the adversary’s available techniques, we classify the following attacks:
Blind attack: The adversary has no prior knowledge of the legitimate user’s eye movement and brainwave patterns. To execute the attack, the attacker wears the user’s XR headset and attempts authentication with their own biometric samples. However, since the attacker cannot gain any advantage from observing the subject during authentication—due to brainwave data being completely resistant to observation and eye movement data requiring specialized devices—this attack is effectively equivalent to a mimic attack. In the mimic attack scenario, other threat models consider observers without any additional capabilities (unaided eye) (Zhang et al., 2024).
Random input attack: We consider an adversary capable of circumventing the XR interface to gain access to the API of our biometric system, enabling them to input arbitrary feature vectors. The adversary’s goal is to find a feature vector that is close to the genuine user’s feature vector. Following Zhao et al. (Zhao et al., 2020), we assume that the feature vectors are normalized and that the number of features is publicly known, with values between 0 and 1. To implement the attack, we generate 1 million samples with 32 values randomly selected from a uniform distribution between 0 and 1. Then, we compare them with the normalized feature vector of the legitimate users.